The gravity on the moon is 1.63 m/s^2. If a ball with a mass of 5 grams is dropped on the moon from a height of 1.5m . Determine the time for the ball to fall on the surface of the moon


1
Expert's answer
2021-05-25T16:48:14-0400

Gives


Gravity of moon

g=a=1.63m/sec2

Mass(M)=5g="5\times10^{-3}kg"

Hight(h)=1.5m

Newton's 2nd law

"S=ut+\frac{1}{2}at^2"

"u=0m\/sec"

"S=0+\frac{1}{2}at^2"

"t=\sqrt\frac{2S}{a}"

Put value


t"=\sqrt\frac{2\times1.5}{1.63}"

t=1.36 sec